# Lite-Mind: Towards Efficient and Versatile Brain Representation Network

### 06 Dec 2023 v1

@@ A.1 Implementation details
Hyper-parameters. On the NSD dataset, during training DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, $\alpha=1$, and $\beta=0$, and CLIP’s contrastive loss is unidirectional for image retrieval and fMRI retrieval. Owning to the lightweight nature of Lite-Mind, the batch size is set to 1500, and the learning rate is 1.16e-3. For Subject 1, 2, 5, 7, Pyramid is set to [512, 256, 128, 257] respectively, and the corresponding number of layers and the patch size are set to [2, 10, 2, 4] and [480, 440, 460, 450], respectively.
About LAION-5B retrieval, for DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, while the weight decay is 6.02e-2 for diffusion projector. CLIP’s contrastive loss is unidirectional for image retrieval, $\alpha=0.5$, and $\beta=0.5$, the batch size is set to 80 and the learning rate is 1.16e-3 for DFT Backbone while 1.1e-4 for diffusion projector.
@@ B.1 Additional Results of Retrieval
To demonstrate the applicability of our method, we also conducted experiments on three other subjects, i.e., Subject 2, Subject 5, and Subject 7, on the NSD dataset, and the experimental results are shown in Table 3. In order to control model size similarity, we did not make more targeted model adjustments for subjects with shorter voxel lengths, but instead used smaller patches. It can be seen that the retrieval accuracies of Subjects 1 and 2 are greater than 90, and the retrieval accuracies of Subjects 5 and 7 are also above 75, The results prove that DFT Backbone can efficiently work on different subjects. Note that MindEye only includes the results of the overall model in Subjects 2, 5, and 7 in its appendix, and it does not evaluate the effect of individual MLP Backbone. As a result, Table 3 does not present any evaluation results for MindEye.
@@
Lite-Mind(Subj 1) 15724 12.51M 94.3% 97.4%
Lite-Mind(Subj 2) 14278 12.49M 93.4% 98.2%
@@
Lite-Mind(Subj 7) 12682 12.47M 75.0% 82.3%
@@ B.2 Additional Results of LAION-5B Retrieval
In order to better reflect the retrieval performance of Lite-Mind on LAION-5B, we presented the performance indicators of LAION-5B retrieval substitution reconstruction for other subjects, i.e., Subject 2, Subject 5, and Subject 7, in Table 4, corresponding to the average performance of subjects in Table 0(b). Similarly, the visualization results of the other subjects in Figure 7 correspond to the image samples of Subject 1 in Figure 4. Based on the comprehensive table and graph, it can be found that Lite-Mind has good generalization on all four subjects, verifying the LAION-5B retrieval ability of Lite-Mind on different subjects. Meanwhile, as shown in Figure 7, the retrieval performance of LAION-5B completely depends on the retrieval accuracy of the CLS model in the test set, such that images retrieved incorrectly in the test set may also have retrieval bias on LAION-5B, for example treating a teddy bear as an image of a cat or dog as shown in Figure 7. However, both MindEye and Lite-Mind exhibit relatively low retrieval accuracy with aligned CLS token models. In the future, it would be beneficial to explore models that improve the alignment of CLS tokens or employ more efficient methods to directly perform retrieval through hidden layers in LAION-5B.
@@ B.3 Additional Results of Zero-shot Classification
B.4 Additional Ablations
In this section, we delved into the dependence of Lite-Mind on data volume and feature length on the NSD dataset and explored the specific impact of different cerebral cortexes on retrieval accuracy, not just nsdgenal. All the experimental results on the NSD dataset are derived from Subject 1, with a retrieving pool size of 300.
B.4.1 Retrieval with varying dataset size
In order to better investigate the versatile representation of DFT Backbone, we conducted data volume experiments on the NSD dataset, as shown in Table 6. Among them, All data corresponds to all trials of every single Subject on the NSD dataset, and 24890 fMRI-image pairs are used for training; Half data refers to randomly removing half of the training data; 2-session refers to using only 1500 randomly selected fMRI-image pairs for training, which corresponds to the two scan sessions of the fMRI experiment. In the table, we report the results of MLP Backbone and diffusion models for MindEye. The experimental results indicate that even in the same NSD dataset, the dependence of Lite-Mind on the number of training samples is smaller than that of MindEye. This also proves that the lightweight model has a wider applicability, i.e., versatility.
Method Parameters Retrieval
@@ B.4.1 Retrieval with varying dataset size
All Data(MindEye) 3B 97.2% 94.7%
Half Data(MindEye) 3B 77.5% 60.8%
2-Sessions(MindEye) 3B 17.9% 12.0%
All Data(Lite-Mind) 12.5M 94.3% 97.4%
Half Data(Lite-Mind) 12.5M 83.4% 91.9%
2-Sessions(Lite-Mind) 12.5M 11.9% 14.1%
Table 6: Retrieval performance with varying dataset size for Subject 1 on the NSD dataset.
@@
B.4.2 Retrieval with varying representation sizes
We conducted additional experiments to verify the impact of representation dimensions on retrieval accuracy, as shown in Table 7. Among them, two different CLIP models (i.e., CLIP ViT/B-32 and CLIP ViT/L-14) were used to extract image representation, with a total of four representation dimensions, by CLS token and the last hidden layer, respectively. From the results, it can be seen that the retrieval accuracy of Lite-Mind is higher when the representation dimension of the image is longer and the representation of the image is richer. The results verify that the fine-grained alignment we mentioned in the main text comes from the rich representation of the image, rather than the fMRI voxel value fully connected heavy MLP Backbone.
Method Representation Parameters Retrieval
[…]
ViT-B/32 CLS 512 6.7M 57.5% 61.8%
ViT-L/14 CLS 768 6.7M 60.3% 64.2%
ViT-B/32 Hidden 50×512 10.4M 91.1% 96.5%
ViT-L/14 Hidden 257×768 12.5M 94.3% 97.4%
Table 7: Retrieval performance with different CLIP representation for Subject 1 on the NSD dataset.
B.4.3 Cerebral cortex
In order to explore the impact of different cortical regions on retrieval accuracy, we used Takagi’s method to extract the stream region of the NSD dataset, covering the visual cortex of nsdgenal, and separately trained the retrieval model for each cortex (early, lateral, parietal, ventral). As shown in Table 8, it can be observed that the early visual cortex has the greatest impact on retrieval accuracy, although with only 5917 voxels Lite-Mind can still achieve retrieval accuracy of 85.0% and 93.4%, which is consistent with Takagi’s research. They believe that visual stimuli are almost dominated by the early visual cortex and indicate that nsdgenal still has a low signal-to-noise ratio, proving that fully connected backbones are unnecessary. Interestingly, for other cortical regions, both lateral and ventral have a good impact on retrieval accuracy, while parietal shows little impact. The findings indicate the necessity of further research on whether the accuracy improvement is related to the longer voxel length of these two regions.
Region Voxel Length Parameters Retrieval
[…]
early 5917 12.4M 85.0% 93.4%
lateral 7799 12.4M 26.9% 29.0%
parietal 3548 12.3M 9.6% 11.2%
ventral 7604 12.3M 18.1% 22.4%
@@ C.2 Retrieval Heatmap
C.3 Visualization in Frequency Domain
We visualized the weights of different patches in different layers of filters, as shown in Figure 10 and 11. Visualization is divided into the real part (left) and imaginary part (right) of filter weights. It can be observed that different channels have varying degrees of attention to different patches, and the frequency domain better captures this characteristic. Interestingly, the weight of the imaginary part for the first and last patches is almost always 0, indicating that the noise is distributed in these two patches.
(a) The real part and imaginary of patch weights for CLS token channel.
(b) The real part and imaginary of patch weights for 140th channel.
Figure 10: Weights visualization for DFT Backbone of $257\times 768$ representation length.
(a) The real part and imaginary of patch weights for the 1st channel.
(b) The real part and imaginary of patch weights for the 60th channel.
Figure 11: Weights visualization for DFT Backbone of 768 representation length.
C.4 Information Alignment
@@
Appendix D Efficiency Discussion
D.1 Theoretical Analysis.
Theorem. Suppose that H is the representation of raw fMRI voxel patches and $\mathcal{H}$ is the corresponding frequency components of the spectrum, then the energy of a voxel patch series in the spatial domain is equal to the energy of its representation in the frequency domain. Formally, we can express this with the above notations:
@@ Appendix D Efficiency Discussion
Where $\mathcal{H}(f)=\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v$, $v$ is the spatial/channel dimension, $f$ is the frequency dimension.
@@
Given the representation of raw voxel patch series $H\in\mathbb{R}^{C\times N}$, let us consider performing integration in either the $C$ dimension (channel dimension) or the $N$ dimension (spatial dimension), denoted as the integral over $v$, then
@@
where $\textbf{H}^{*}(v)$ is the conjugate of $\textbf{H}(v)$. According to IDFT, $\textbf{H}^{*}(v)=\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f$, we can obtain
$\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v$ $=\int_{-\infty}^{\infty}{\textbf{H}(v)}[\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f]\,\mathrm{d}v$ (9)
=*$=\int_{-\infty}^{\infty}{\mathcal{H}}^{*}(f)[\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v]\,\mathrm{d}f$
=*$=\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}{\mathcal{H}(f)}\,\mathrm{d}f$
$=\int_{-\infty}^{\infty}{|\mathcal{H}(f)|}^{2}\,\mathrm{d}f$
@@
Therefore, the energy of a voxel patch series in the spatial domain is equal to the energy of its representation in the frequency domain.
Complexity Analysis. For a fMRI voxel with a length of $m$, we divide it into $n$ $l$-length patches ($m=nl$) with $C$ channel numbers. Assuming $L_{1}$ and $L_{2}$ are the layer depths of MLP and DFT respectively, the middle layer dimension of MLP is $D_{1}$, and the alignment representation dimension is $N\times D_{2}$, where $N$ is the number of the last hidden layer of CLIP. The time complexity of MLP Backbone is $O(mD_{1}+L_{1}D_{1}^{2}+ND_{1}D_{2})$. For DFT Backobone, the time complexity of patchfy is $O(mC)$, and the time complexity of DFT, IDFT, and filtering for each layer is $O(2nC\operatorname{log}n+nC)$. The time complexity of each layer’s tiny MLP is $O(8nC^{2})$. Thus the time complexity of the entire DFT Backbone is:
$O(mC+L_{2}(2nC\operatorname{log}n+nC+8nC^{2})+nD_{2})$ (10)
$=O(mC+nC\operatorname{log}nL_{2}+nC^{2}L_{2}+nD_{2})$
Quantitative analysis algorithm complexity for DFT Backbone is shown in Table 9.
Model FLOPs(G) Parameters Image Retrieval Brain Retrieval
MindEye 5.66 996M 88.8% 84.9%
Pyramid-18 0.80 12.5M 94.3% 97.4%
DFT-18 0.63 9.9M 91.8% 97.3%
DFT-12 0.43 6.6M 91.0% 97.2%
DFT-6 0.22 3.4M 90.6% 97.2%
DFT-1 0.05 0.7M 83.5% 94.0%
DFT-0 0.02 0.2M 65.8% 74.7%
Table 9: Quantitative complexity analysis for DFT Backbone with different layers and MLP Backbone+Projector for MindEye.

### 12 Mar 2024 v2

@@ A.1 Implementation details
Hyper-parameters. On the NSD dataset, during training DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, $\alpha=1$, and CLIP’s contrastive loss is unidirectional for image retrieval and fMRI retrieval. Owning to the lightweight nature of Lite-Mind, the batch size is set to 500, the learning rate is 1e-3, patch size is set to 480, and Filter library size is 4. Filter block layers are the same for all subjects.
About LAION-5B retrieval, for DFT Backbone, the weight decay is set to 7, $\tau$ is $1/e^{8}$, while the weight decay is 6.02e-2 for diffusion projector. CLIP’s contrastive loss is unidirectional for image retrieval, $\alpha=0.5$, the batch size is set to 80 and the learning rate is 1e-3 for DFT Backbone while 1.1e-4 for diffusion projector. Note that hyper-parameters in Diffusion Projector are the value of the open-source DALLE·2 from https://github.com/lucidrains/DALLE2-pytorch.
@@ B.1 Additional Results of Retrieval
To demonstrate the applicability of our method, we also conducted experiments on three other Subjects, i.e., Subject 2, Subject 5, and Subject 7, on the NSD dataset, and the experimental results are shown in Table 7. In order to control model size similarity, we did not make more targeted model adjustments for subjects with shorter voxel lengths. It can be seen that the retrieval accuracies of Subjects 1 and 2 are greater than 90%, and the retrieval accuracies of Subjects 5 and 7 are also above 80%. The results prove that DFT Backbone can efficiently work on different subjects. Note that MindEye only includes the results of the overall model in Subjects 2, 5, and 7 in its Appendix, and it does not evaluate the effect of individual MLP Backbone. As a result, Table 7 does not present any evaluation results for MindEye.
@@
Lite-Mind(Subj 1) 15724 12.51M 94.6% 97.4%
Lite-Mind(Subj 2) 14278 12.49M 94.1% 98.2%
@@
Lite-Mind(Subj 7) 12682 12.47M 81.7% 82.3%
@@ B.2 Additional Results of LAION-5B Retrieval
In order to better reflect the retrieval performance of Lite-Mind on LAION-5B, we presented the performance indicators of LAION-5B retrieval substitution reconstruction for other subjects, i.e., Subject 2, Subject 5, and Subject 7, in Table 8, corresponding to the average performance of subjects in Table 0(b). Similarly, the visualization results of the other subjects in Figure 7 correspond to the image samples of Subject 1 in Figure 4. Based on the comprehensive table and graph, it can be found that Lite-Mind has good generalization on all four subjects, verifying the LAION-5B retrieval ability of Lite-Mind on different subjects. Meanwhile, as shown in Figure 7, the retrieval performance of LAION-5B completely depends on the retrieval accuracy of the CLS model in the test set, such that images retrieved incorrectly in the test set may also have retrieval bias on LAION-5B, for example treating a teddy bear as an image of a cat or dog as shown in Figure 7. However, both MindEye and Lite-Mind exhibit relatively low retrieval accuracy with aligned CLS embedding models. In the future, it would be beneficial to explore models that improve the alignment of CLS embeddings or employ more efficient methods to directly perform retrieval through hidden layers in LAION-5B.
@@ B.3 Additional Results of Zero-shot Classification
B.4 Hyper-parameters Experiment
In this section, we explore the influence of some hyperparameters on the model, including patch size and number of filters $M$, to verify the parameter sensitivity of the model. All the experimental results on the NSD dataset are from Subject 1, with a retrieving pool size of 300.
Patch size. We conducted experiments by varying the patch size as presented in Table 10. The results exhibit stable retrieval accuracy above 90%, which indicates that our DFT Backbone is not sensitive to part size. Since we found patch size has less effect on retrieval accuracy, we chose a relatively large patch size (e.g., 480) to ensure fewer parameters and faster convergence.
Patch size Parameters Retrieval
@@ B.4 Hyper-parameters Experiment
50 14.0M 91.6% 96.4%
200 12.7M 92.9% 96.9%
480 12.5M 94.6% 97.1%
600 12.3M 94.2% 97.0%
900 11.9M 90.1% 95.5%
Table 10: Retrieval performance for different patch size on the NSD dataset.
Filter library size. We conducted experiments by varying the Filter library size as presented in Table 11. Specifically, there is a significant performance improvement from $M=1$ to $M=4$, while a slight fluctuation is observed for $M=4/8$. By setting $M=4$, the model has the ability to acquire diverse and distinct feature patterns from various dimensions of the frequency response while still maintaining an appropriate computational cost. Therefore, we determine that $M=4$ is the optimal choice on the NSD dataset.
Filters Parameters Retrieval
[…]
1 9.0M 89.8% 96.4%
2 10.2M 93.5% 97.4%
4 12.5M 94.6% 97.1%
8 17.2M 94.4% 96.7%
Table 11: Retrieval performance for different Filter library size on the NSD dataset.
@@ C.2 Retrieval Heatmap
C.3 Information Alignment
@@
C.4 More Retrieval Cases
In this section, we visualize retrieval failure cases in Figure 10, although only dozens of images are not in the Top 1. From the left half of the Figure, it can be seen that though not in Top 1, Lite-Mind still retrieved ground-truth in Top 2, and images in Top 4 are similar(either semantically similar or structurally similar, eg. animals in the wild or an airplane on the runway). On the contrary, a smaller proportion of ground-truth did not appear within the Top 4, as shown in the right half of the Figure. From the perspective of the images themselves, most of the scenes are too complex, even abstract (as shown in the second image), which may be the reason for the retrieval failure.
Figure 10: Partial failure retrieval results of Lite-Mind on all 982 test images for Subject 1. The number below each image represents the similarity score.
(a) The real part and imaginary of a filter weights.
(b) The real part and imaginary of another filter weights from the same Filter Library.
Figure 11: Weights visualization for DFT Backbone of $257\times 768$ embedding length.
(a) The real part and imaginary of a filter weights.
(b) The real part and imaginary of another filter weights from the same Filter Library.
Figure 12: Weights visualization for DFT Backbone of 768 CLS embedding length.
C.5 Visualization in Frequency Domain
We visualized the weights of Filter library in Filter Blocks, as shown in Figure 11 and 12. Visualization is divided into the real part (left) and imaginary part (right) of filter weights. It can be observed that different filters have varying degrees of attention to different tokens, and the frequency domain better captures this characteristic. Interestingly, the weight of the imaginary part for the first and last tokens is almost always 0, indicating that the noise is distributed in these two tokens.
Appendix D Theoretical Analysis.
D.1 Complex Multiplication.
For two complex number values $\mathcal{Z}_{1}=(a+jb)$ and $\mathcal{Z}_{2}=(c+jd)$, where $a$ and $c$ is the real part of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ respectively, $b$ and $d$ is the imaginary part of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ respectively. Then the multiplication of $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ is calculated by:
$\mathcal{Z}_{1}\mathcal{Z}_{2}$ $=(a+jb)(c+jd)$ (10)
$=ac+j^{2}bd+jad+jbc$
$=(ac-bd)+j(ad+bc)$
D.2 Theorem Proof.
Theorem 1. Suppose that H is the representation of raw fMRI voxel tokens and $\mathcal{H}$ is the corresponding frequency components of the spectrum, then the energy of voxel tokens in the spatial domain is equal to the energy of its representation in the frequency domain. Formally, we can express this with the above notations:
@@ C.5 Visualization in Frequency Domain
Where $\mathcal{H}(f)=\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v$, $v$ is the token dimension, $f$ is the frequency dimension.
@@
Given the representation of raw voxel token series $H\in\mathbb{R}^{C\times N}$, let us consider performing integration in the $N$ dimension (spatial dimension), denoted as the integral over $v$, then
@@
where $\textbf{H}^{*}(v)$ is the conjugate of $\textbf{H}(v)$. According to IDFT, $\textbf{H}^{*}(v)=\int_{-\infty}^{\infty}{\mathcal{H}^{*}\textbf{}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f$, we can obtain
$\int_{-\infty}^{\infty}{|\textbf{H}(v)|}^{2}\,\mathrm{d}v&=\int_{-\infty}^{\infty}{\textbf{H}(v)}[\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}e^{-j2{\pi}fv}\,\mathrm{d}f]\,\mathrm{d}v\\ &=\int_{-\infty}^{\infty}{\mathcal{H}}^{*}(f)[\int_{-\infty}^{\infty}{|\textbf{H}(v)|}e^{-j2{\pi}fv}\,\mathrm{d}v]\,\mathrm{d}f\\ &=\int_{-\infty}^{\infty}{\mathcal{H}^{*}(f)}{\mathcal{H}(f)}\,\mathrm{d}f\\ &=\int_{-\infty}^{\infty}{|\mathcal{H}(f)|}^{2}\,\mathrm{d}f\end{aligned}$
@@
Therefore, the energy of a voxel token series in the spatial domain is equal to the energy of its representation in the frequency domain.
Theorem 2. Given the voxel token series input H and its corresponding frequency domain conversion H, the operations of frequency-domain MLP on H can be represented as global convolutions on H in the spatial domain. This can be given by:
$\mathcal{H}\mathcal{W}+\mathcal{B}=\mathcal{F}(\textbf{H}*W+B)$ (14)
Where $\mathcal{F}$ is DFT, *** is a circular convolution, $\mathcal{W}$ and $\mathcal{B}$ are the complex number weight and bias, and $W$ and $B$ are the weight and bias in the spatial domain.
Proof.
Suppose that we conduct operations in the N (i.e., token dimension), then
$\mathcal{F}(\textbf{H}(v)*W(v))=\int_{-\infty}^{\infty}{(\textbf{H}(v)*W(v))e^{-j2\pi fv}}\,\mathrm{d}v$ (15)
According to convolution theorem, $\textbf{H}(v)*W(v)=\int_{-\infty}^{\infty}(\textbf{H}(\tau)W(v-\tau))\,\mathrm{d}\tau$, then
$\mathcal{F}(\textbf{H}(v)*W(v))&=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{(\textbf{H}(\tau)W(v-\tau))e^{-j2\pi fv}}\,\mathrm{d}\tau}\,\mathrm{d}v\\ &=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(v-\tau)e^{-j2\pi fv}}\,\mathrm{d}v}\textbf{H}(\tau)\,\mathrm{d}\tau\end{aligned}$
Let $x=v-\tau$,then
$\mathcal{F}(\textbf{H}(v)*W(v))&=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(x)e^{-j2\pi f(x+\tau)}}\,\mathrm{d}x}\textbf{H}(\tau)\,\mathrm{d}\tau\\ &=\int_{-\infty}^{\infty}{\int_{-\infty}^{\infty}{W(x)e^{-j2\pi fx}e^{-j2\pi f\tau}}\,\mathrm{d}x}\textbf{H}(\tau)\,\mathrm{d}\tau\\ &=\int_{-\infty}^{\infty}{(\textbf{H}(\tau)*)e^{-j2\pi f\tau}}\,\mathrm{d}\tau\int_{-\infty}^{\infty}{(\textbf{W}(x)*)e^{-j2\pi fx}}\,\mathrm{d}x\\ &=\mathcal{H}(f)\mathcal{W}(f)\end{aligned}$
Accordingly, $\textbf{H}(v)*W(v)$ in the spatial domain is equal to $\mathcal{H}(f)\mathcal{W}(f)$ in the frequency domain. Therefore, the operations of FreMLP ($\mathcal{H}\mathcal{W}+\mathcal{B}$) in the token dimension (i.e., v = N ) are equal to the operations ($\mathcal{F}(\textbf{H}*W+B)$) in the spatial domain. This implies that frequency-domain MLPs can be viewed as global convolutions in the spatial domain. Proved. ∎
D.3 Complexity Analysis
For a fMRI voxel with a length of $l$, we divide it into $n$ patches. Assuming $L_{1}$ and $L_{2}$ are the layer depths of MLP Backbone and DFT Backbone respectively, the middle layer dimension of MLP Backbone is $D$, and the alignment embedding dimension is $n^{\prime}\times D^{\prime}$, where $n^{\prime}$ is the the number of tokens of CLIP. The time complexity of MLP Backbone is $O(lD+L_{1}D^{2}+n^{\prime}DD^{\prime})$. For DFT Backobone, the time complexity of patchify and tokenization is $O(lD^{\prime})$, and the time complexity of DFT, IDFT, and filtering for each layer is $O(2nD^{\prime}\operatorname{log}n+nD^{\prime})$. The time complexity of FreMLP is $O(2nD^{\prime}\operatorname{log}n+2nn^{\prime}D^{\prime}+2n^{\prime}D^{\prime})$. Thus the time complexity of the entire DFT Backbone is:
$O(4nD^{\prime}\operatorname{log}n+(n+2nn^{\prime}+2n^{\prime})D^{\prime})$ (18)
$=O((n\operatorname{log}n+nn^{\prime}+n^{\prime})D^{\prime})$